# Residual subspace evolution strategies for nonlinear inverse problems

### v1

@@
Abstract
Nonlinear inverse problems often feature noisy, non-differentiable, or expensive residual evaluations that make Jacobian-based solvers unreliable. Popular derivative-free optimizers such as natural evolution strategies (NES) or Powell’s NEWUOA still assume smoothness or expend many evaluations to maintain stability. Ensemble Kalman inversion (EKI) relies on empirical covariances that require preconditioning and scale poorly with residual dimension.
We introduce residual subspace evolution strategies (RSES), a derivative-free solver that samples Gaussian probes around the current iterate, builds a residual-only surrogate from their differences, and recombines the probes through a least-squares solve yielding an optimal update without forming Jacobians or covariances. Each iteration costs $k+1$ residual evaluations, where $k\ll n$ for $n$-dimensional problems, with $O(k^{3})$ linear algebra overhead.
Benchmarks on calibration, regression, and deconvolution problems demonstrate consistent misfit reduction in both deterministic and stochastic settings. RSES matches or surpasses xNES and NEWUOA while staying competitive with EKI under matched evaluation budgets, particularly when smoothness or covariance assumptions fail.
Keywords residual subspace evolution strategies, nonlinear inverse problems, derivative-free optimisation, ensemble methods, Tikhonov regularisation, stochastic search
@@ 1 Introduction
Many estimation, calibration, and imaging tasks reduce to nonlinear inverse problems described by the residual map in Eq. 1.
$F:\mathbb{R}^{n}\to\mathbb{R}^{m},\qquad F(\mathbf{x})\approx 0.$ (1)
Classical solvers such as Gauss-Newton, sequential quadratic programming, and trust-region variants perform well when derivatives are stable and forward maps are clean, but they falter once residuals are noisy, non-differentiable, or expensive to query. Derivative-free quadratic modelers like NEWUOA [powell2006newuoa] relax derivative access yet still assume smoothness and spend many evaluations to maintain reliable local surrogates.
@@
Population methods build robustness from sample statistics. Ensemble Kalman inversion [evensen2009data] regularises steps with empirical covariances, and natural evolution strategies such as xNES [wierstra2014natural] adapt search directions globally. They usually need sizeable populations to keep their estimates stable, which inflates cost when each residual call is expensive.
@@
Residual subspace techniques point to a lighter alternative. Anderson acceleration [anderson_iterative_1965, walker2011anderson] and direct inversion in the iterative subspace (DIIS) [pulay1982] recycle residual differences as proxies for Jacobian actions inside small affine subspaces, often converging rapidly without explicit derivatives. While these methods target fixed-point problems rather than general optimization and do not use ensembles, they demonstrate that residual evaluations alone can recover low-rank Jacobian information.
@@
Building on this insight, residual subspace evolution strategies (RSES) blend these ideas for noisy inverse problems. Each iteration perturbs the previous iterate with $k$ Gaussian draws, stacks the resulting residual differences, and solves a least-squares problem on the probed subspace. The resulting coefficients recombine the probes into an optimal update without forming Jacobians.
@@
Our contributions are threefold: (i) we derive RSES as the solution of a minimal-residual problem on the probe subspace, (ii) we provide parameter guidance for the ridge term, probe count, and probe scale, and (iii) we report deterministic and Monte Carlo benchmarks on systems ranging from two-parameter calibration to $\sim 4000$-dimensional MLP regression tasks. Under matched evaluation budgets, RSES reduces residuals more effectively than derivative-free and gradient-based baselines when their model or smoothness assumptions break.
Section 2 develops the probe-based surrogate and ridge recombination, Section 3 details algorithmic choices, Section 4 reports experiments, and Section 5 summarises implications and future work.
2 Mathematical derivation
@@
The RSES update solves a regularised minimal-residual problem on a span of Gaussian probes. Each iteration samples Gaussian perturbations and weights them so that their recombination forms the update direction. We write vectors and matrices in bold: $\mathbf{x}_{t}$ denotes the current iterate, $\mathbf{r}_{t}\coloneqq F(\mathbf{x}_{t})$ its residual, $\mathbf{P}$ stacks the probes $\mathbf{p}_{i}$, $\mathbf{B}$ collects their residual differences, and $\mathbf{w}$ stores the recombination weights. We denote by $\epsilon$ a small numerical floor, default $10^{-8}$, that stabilises the ridge update. The derivation proceeds in three steps: (1) build a linear surrogate of the residuals on the probe span anchored at the current iterate, (2) pick coefficients from a least-squares solve, and (3) turn those coefficients into a state update that stays inside the probed space.
Step 1: Probe surrogate.
We frame the task through the residual energy:
@@
At iteration $t$ we draw $k$ Gaussian perturbations with scale $\sigma$ to populate the probe and residual matrices
@@
with each probe built from a scaled standard normal draw $\mathbf{g}_{i}\sim\mathcal{N}(0,I_{n})$ so that $\mathbf{p}_{i}=\sigma\mathbf{g}_{i}$ and $\mathbf{r}_{t}=F(\mathbf{x}_{t})$. A first-order expansion of $F$ around $\mathbf{x}_{t}$ with Jacobian $\mathbf{J}_{t}$ yields
@@
so, for any coefficient vector $\mathbf{w}$,
@@
This surrogate reuses measured residual differences and keeps the expansion centred at $\mathbf{x}_{t}$ without extra constraints.
Step 2: Ridge recombination.
We minimise $\phi(\mathbf{x}_{t}+\mathbf{P}\mathbf{w})$ under the surrogate (6) with a small Tikhonov stabiliser:
@@
The ridge term keeps the Gram matrix well behaved; with $\lambda>0$, $\mathbf{B}^{\top}\mathbf{B}+\lambda I_{k}$ stays positive definite, and the normal equations reduce to the $k\times k$ solve
@@
which produces coefficients that balance residual reduction with coefficient magnitude.
Step 3: State update.
Applying the recovered weights to the probes yields
@@
The step lives in the probe span, stays centred at $\mathbf{x}_{t}$, and mirrors a Gauss-Newton projection on the surrogate rather than on the full space.
In the idealised linear regime on the probe span, the ridge solve produces a Gauss-Newton projection that does not raise the residual. Lemma 1 in Appendix A shows that the exact surrogate yields $\phi(\mathbf{x}_{t+1})\leq\phi(\mathbf{x}_{t})$, with strict descent when the projected gradient on the probe span is nonzero.
3 Algorithm and parameter selection
The surrogate in Section 2 becomes a compact loop that keeps each move inside the span of the current probes. At each iteration we evaluate the residual, draw $k$ Gaussian probes, record how the residual changes along them, solve a ridge-stabilised Gram system, and recombine the probes into a step. These pieces and the parameter rules below produce Algorithm 1.
The update mirrors a regularised Gauss-Newton step on the probed subspace. The ridge term regularises the recombination coefficients and guards against nearly singular systems; we floor $\lambda$ at $\epsilon$ to keep the solve well conditioned.
Parameter selection
The loop exposes three user choices: the probe count $k$, the probe scale $\sigma$, and the ridge scale $\beta$. We tie the ridge parameter to the residual norm through
Probe count.
With an evaluation budget of $N_{\text{eval}}$ residual calls, we prescribe
which keeps the linear solve small while widening the span slowly with $m$. This prescription lacks deep theoretical justification; we found it adequate across our benchmarks. Increase the coefficient on $\log m$ when the residual shows strong anisotropy and the Gram matrix stays well conditioned. The logarithmic scaling loosely aligns with the Johnson–Lindenstrauss lemma if one views the probe subspace as approximating an $m$-dimensional residual space through random projection.
Probe scale.
The probe amplitude should match the problem units. For unit-scaled variables, we recommend $\sigma=0.05$. For unscaled problems, set $\sigma$ to a modest fraction (e.g., 0.010.010.10.1) of the prior standard deviation or the norm of a trusted starting point. If the first steps stall, double $\sigma$; if residuals spike, halve it.
Ridge scale.
The ridge keeps the Gram matrix stable without dominating the residual term. We set $\beta=10^{-5}$, which makes $\lambda_{t}$ decay with $\lVert\mathbf{r}_{t}\rVert_{2}^{2}$ as the iterate improves. Larger $\beta$ damps oscillations at the cost of smaller steps; smaller values accelerate progress when the probes align well but may destabilize the solve. For severely ill-conditioned problems, increase $\beta$ to $10^{-4}$ or $10^{-3}$.
Figure 1 sweeps these parameters on the nonlinear deconvolution benchmark. Moderate changes leave performance stable; very small probe counts or scales slow progress, while excessively large ridge values yield overly conservative steps.
@@ 3 Algorithm and parameter selection
Computational complexity.
Each iteration requires $k+1$ evaluations of $F$, forms $\mathbf{B}^{\top}\mathbf{B}$ and $\mathbf{B}^{\top}\mathbf{r}$ in $O(mk^{2})$ time, solves a $k\times k$ system in $O(k^{3})$ time, and computes $\mathbf{P}\mathbf{w}$ in $O(nk)$ time. Storage for $\mathbf{P}$ and $\mathbf{B}$ costs $O(nk+mk)$ memory, while the Gram factorization adds $O(k^{2})$, yielding total per-iteration memory of $O(nk+mk+k^{2})$. The per-iteration computational cost is $O((k+1)C_{F}+mk^{2}+k^{3}+nk)$ for residual cost $C_{F}$. When residual evaluations dominate, the linear algebra overhead remains negligible.
@@ 4 Numerical experiments
We use the parameter choices in Section 3. Each solver receives 7,5007{,}500 residual evaluations and a maximum runtime of $30$ seconds. We record the best-so-far residual norm, the corresponding parameter error $\lVert\mathbf{x}-\mathbf{x}_{\star}\rVert_{2}$, and cumulative runtime. RSES sets $k$ from Eq. 11 and maintains a fixed probe scale $\sigma$.
We average ten independent trials per benchmark and compare RSES against NEWUOA [powell2006newuoa], xNES [wierstra2014natural], Gauss-Newton (GN) with finite-difference Jacobians, and ensemble Kalman inversion (EKI) [iglesias2013ensemble]. Tables report mean residual, parameter error, and elapsed time at a shared stopping index: the first evaluation where the mean RSES distance falls within $1\%$ of its minimum. Every solver reports its values at this index, enabling fair comparison at a matched stopping point rather than arbitrary terminal iterates. Curves show mean best-so-far distance with root-mean-square bands on a shared evaluation axis.
The benchmarks escalate in difficulty, covering: (i) a $2\times 2$ linear system, (ii) low-dimensional stochastic Brownian calibration, (iii) noisy multilayer-perceptron regression with up to 4,3534{,}353 parameters, (iv) nonlinear deconvolution with $128$ parameters, and (v) the same deconvolution with perturbed residual weighting.
@@ 4.1 Linear algebraic system
We first solve the $2\times 2$ linear system
@@ 4.2 Brownian drift and diffusion
and we tune the parameters gathered in
@@
$\hat{m}(\boldsymbol{\theta})=\tfrac{1}{N}\sum_{j=1}^{N}X_{1}^{(j)},\qquad\hat{v}(\boldsymbol{\theta})=\tfrac{1}{N}\sum_{j=1}^{N}\bigl(X_{1}^{(j)}-\hat{m}(\boldsymbol{\theta})\bigr)^{2},$ (16)
@@
We evolve each trajectory with Euler using $32$ steps of size $1/32$ and average 4,0964{,}096 paths. RSES tracks the target despite Monte Carlo noise and achieves the smallest residual and parameter error. EKI follows closely in accuracy but runs slower. NEWUOA converges fastest yet stops far from the target. xNES reaches moderate accuracy with cost similar to RSES. GN overshoots because noisy finite differences distort the Jacobian approximation (Tables 2 and 3).
@@ 4.3 Multilayer perceptron regression
We add a one-dimensional regression task to probe a nonconvex landscape. The target signal follows Eq. 18 with additive Gaussian noise of standard deviation $0.05$.
$y(x)=\sin(3x)+0.3x.$ (18)
We fit a two-hidden-layer multilayer perceptron with widths $8$, $16$, $32$, and $64$ using the smooth activation in Eq. 19 and a Tikhonov penalty of $10^{-6}$.
$\phi(s)=\frac{s}{\sqrt{1+s^{2}}}.$ (19)
This inverse-square-root activation belongs to the softsign family and keeps derivatives bounded [glorot2010understanding]. With scalar input $x$, the network reads
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{1}\times 1}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{2}\times d_{1}}$, $\mathbf{W}_{3}\in\mathbb{R}^{1\times d_{2}}$, and $(d_{1},d_{2})$ is one of the four width pairs above. We define the Tikhonov residual
$F(\boldsymbol{\theta})=\begin{bmatrix}\hat{y}_{\boldsymbol{\theta}}(x)-y\\ \sqrt{\lambda/2}\,\boldsymbol{\theta}\end{bmatrix}$, (21)
which stacks data misfit and $\ell_{2}$ regularisation. We measure loss as the squared residual norm,
@@
This loss surface contains broad plateaus and narrow valleys because the activation saturates and the regulariser couples all parameters, so gradient methods can stall even on this small dataset.
RSES uses $k=20$ (following $k=4+\lfloor 3\log m\rfloor$ with $m=256$ data points; the Tikhonov terms do not contribute to $m$ since the regularisation does not add problem complexity), $\sigma=0.01$, ridge scale $\beta=10^{-5}$, and approximately 20,00020{,}000 residual evaluations spread across $952$ iterations. Each iteration evaluates Gaussian probes of the Tikhonov residual without forming gradients. Adam [kingma2014adam] receives the same 20,00020{,}000 gradient evaluations. We omit xNES, EKI, and GN on this task because their sampling and covariance costs scale poorly with these parameter counts. NEWUOA uses the same evaluation budget but slows sharply, exceeding the $30$-second time limit once the width reaches $16$. Table 3 reports final losses $\mathcal{L}(\boldsymbol{\theta})$ and wall-clock time. RSES attains the lowest loss at every width while running faster than Adam and orders of magnitude faster than NEWUOA on wider networks. Adam remains competitive in loss but slower in time. NEWUOA stalls at substantially higher loss once the width reaches $16$.
Figure 4 shows residual histories across widths and the resulting best fits per algorithm. RSES descends steadily while Adam plateaus before recovering. The RSES fits track the noisy target closely without overfitting. These results indicate that RSES navigates the nonconvex MLP landscape reliably at this scale, though larger data regimes will require batching and may still favor tuned gradient-based training.

### v2

@@
Abstract: Nonlinear inverse problems pervade engineering and science, yet noisy, non-differentiable, or expensive residual evaluations routinely defeat Jacobian-based solvers. Derivative-free alternatives either demand smoothness, require large populations to stabilise covariance estimates, or stall on flat regions where gradient information fades.
This paper introduces residual subspace evolution strategies (RSES), a derivative-free solver that draws Gaussian probes around the current iterate, records how residuals change along those directions, and recombines the probes through a least-squares solve to produce an optimal update. The method builds a residual-only surrogate without forming Jacobians or empirical covariances, and each iteration costs just $k+1$ residual evaluations with $O(k^{3})$ linear algebra overhead, where $k$ remains far smaller than the parameter dimension.
Benchmarks on calibration, regression, and deconvolution tasks show that RSES reduces misfit consistently across deterministic and stochastic settings, matching or exceeding xNES, NEWUOA, Adam, and ensemble Kalman inversion under matched evaluation budgets. The gains are most pronounced when smoothness or covariance assumptions break, suggesting that lightweight residual-difference surrogates can reliably guide descent where heavier machinery struggles.
Keywords: residual subspace evolution strategies, nonlinear inverse problems, derivative-free optimisation, ensemble methods, Tikhonov regularisation, stochastic search
@@ 1 Introduction
Many estimation, calibration, and imaging tasks reduce to finding parameter vectors $\mathbf{x}\in\mathbb{R}^{n}$ that drive a residual map
$F:\mathbb{R}^{n}\to\mathbb{R}^{m},\qquad F(\mathbf{x})\approx 0$ (1)
toward zero. Classical solvers such as Gauss-Newton, sequential quadratic programming, and trust-region variants perform well when derivatives remain stable and forward models behave smoothly, yet they falter once residuals become noisy, non-differentiable, or expensive to query. Adaptive gradient methods like Adam [kingma2014adam] handle nonconvex objectives with momentum and per-parameter scaling, yet they can stall on broad plateaus or when gradient noise obscures descent directions.
@@
Derivative-free quadratic modelers like NEWUOA [powell2006newuoa] relax the need for derivative access but still assume smoothness, spending many evaluations to maintain reliable local surrogates. Population methods build robustness from sample statistics: ensemble Kalman inversion [evensen2009data] regularises steps with empirical covariances, and natural evolution strategies such as xNES [wierstra2014natural] adapt search directions globally. These approaches usually require sizeable populations to keep their estimates stable, which inflates cost when each residual call is expensive.
@@
Residual subspace techniques point to a lighter alternative. Anderson acceleration [anderson_iterative_1965, walker2011anderson] and direct inversion in the iterative subspace (DIIS) [pulay1982] recycle residual differences as proxies for Jacobian actions inside small affine subspaces, often converging rapidly without explicit derivatives. Although these methods target fixed-point problems rather than general optimization and do not employ ensembles, they demonstrate that residual evaluations alone can recover low-rank Jacobian information.
@@
Building on this insight, the present work introduces residual subspace evolution strategies (RSES), a derivative-free solver designed for noisy inverse problems. Each iteration perturbs the current iterate with $k$ Gaussian draws, stacks the resulting residual differences, and solves a least-squares problem on the probed subspace. The recovered coefficients recombine the probes into an optimal update without forming Jacobians.
@@
This paper derives RSES as the solution of a minimal-residual problem on the probe subspace, provides parameter guidance for the ridge term, probe count, and probe scale, and reports deterministic and Monte Carlo benchmarks on systems ranging from two-parameter calibration to multilayer-perceptron regression tasks with several thousand parameters. Under matched evaluation budgets, RSES reduces residuals more effectively than derivative-free and gradient-based baselines when their model or smoothness assumptions break. The following sections develop the core methodology, present numerical experiments, and conclude with a discussion of limitations and future directions.
2 Core methodology
@@
The RSES update solves a regularised minimal-residual problem on a span of Gaussian probes. Each iteration samples perturbations around the current iterate, records the corresponding residual changes, and weights the perturbations so that their recombination yields an improved estimate. The notation uses bold symbols for vectors and matrices: $\mathbf{x}_{t}$ denotes the current iterate, $\mathbf{r}_{t}\coloneqq F(\mathbf{x}_{t})$ its residual, $\mathbf{P}$ collects the probe directions $\mathbf{p}_{i}$, $\mathbf{B}$ stores their residual differences, and $\mathbf{w}$ holds the recombination weights. A small numerical floor $\epsilon$, set by default to $10^{-8}$, stabilises the ridge update.
Probe surrogate.
The method frames the inverse problem through the residual energy
@@
At iteration $t$, the algorithm draws $k$ Gaussian perturbations with scale $\sigma$, yielding the probe and residual-difference matrices
@@
with each probe constructed from a scaled standard normal draw $\mathbf{g}_{i}\sim\mathcal{N}(0,I_{n})$ so that $\mathbf{p}_{i}=\sigma\mathbf{g}_{i}$. A first-order expansion of $F$ around $\mathbf{x}_{t}$ with Jacobian $\mathbf{J}_{t}$ produces
@@
from which, for any coefficient vector $\mathbf{w}$, the residual at the displaced point satisfies
@@
This surrogate reuses measured residual differences and keeps the expansion centred at $\mathbf{x}_{t}$ without imposing additional constraints.
Ridge recombination.
The next step minimises $\phi(\mathbf{x}_{t}+\mathbf{P}\mathbf{w})$ under the surrogate (6) with a small Tikhonov stabiliser:
@@
The ridge term keeps the Gram matrix well behaved: with $\lambda>0$, the matrix $\mathbf{B}^{\top}\mathbf{B}+\lambda I_{k}$ remains positive definite, and the normal equations reduce to the $k\times k$ system
@@
which delivers coefficients that balance residual reduction against coefficient magnitude.
State update.
Applying the recovered weights to the probes yields the parameter update
@@
The step lives entirely within the probe span, stays centred at $\mathbf{x}_{t}$, and mirrors a Gauss-Newton projection on the surrogate rather than on the full parameter space. In the idealised linear regime on the probe span, the ridge solve produces a descent direction that does not raise the residual energy. Lemma 1 in appendix A establishes that the exact surrogate yields $\phi(\mathbf{x}_{t+1})\leq\phi(\mathbf{x}_{t})$, with strict descent whenever the projected gradient on the probe span is nonzero.
2.1 Algorithm
The derivation above translates directly into a compact iteration that keeps each move inside the span of the current probes. At every step the algorithm evaluates the residual, draws $k$ Gaussian probes, records how the residual changes along them, solves a ridge-stabilised Gram system, and recombines the probes into a step. Algorithm 1 summarises this procedure.
@@ 2.1 Algorithm
2.2 Parameter selection
The iteration exposes three user choices: the probe count $k$, the probe scale $\sigma$, and the ridge scale $\beta$. A residual-dependent ridge parameter links these through
Probe count.
Given an evaluation budget of $N_{\text{eval}}$ residual calls, a practical prescription sets
which keeps the linear solve small while widening the span slowly with the residual dimension $m$. This rule lacks deep theoretical justification but proved adequate across the benchmarks reported below. Increasing the coefficient on $\log m$ helps when the residual shows strong anisotropy and the Gram matrix stays well conditioned. The logarithmic scaling loosely aligns with the Johnson-Lindenstrauss lemma when one views the probe subspace as approximating an $m$-dimensional residual space through random projection.
Probe scale.
The probe amplitude should match the problem units. For unit-scaled variables, a value of $\sigma=0.05$ works well. For unscaled problems, setting $\sigma$ to a modest fraction of the prior standard deviation or the norm of a trusted starting point typically suffices. If the first steps stall, doubling $\sigma$ often helps; if residuals spike, halving it restores stability.
Ridge scale.
The ridge keeps the Gram matrix stable without dominating the residual term. A default of $\beta=10^{-5}$ makes $\lambda_{t}$ decay with $\lVert\mathbf{r}_{t}\rVert_{2}^{2}$ as the iterate improves. Appendix B shows that this schedule bounds the coefficient norm $\lVert\mathbf{w}\rVert_{2}$ uniformly, independent of the residual matrix $\mathbf{B}$, while maintaining step sizes proportional to the residual near convergence. Larger $\beta$ damps oscillations at the cost of smaller steps; smaller values accelerate progress when the probes align well but may destabilise the solve. For severely ill-conditioned problems, increasing $\beta$ to $10^{-4}$ or $10^{-3}$ restores robustness.
Figure 1 sweeps these parameters on the nonlinear deconvolution benchmark described in section 3.4. Moderate changes leave performance stable; very small probe counts or scales slow progress, while excessively large ridge values yield overly conservative steps.
@@ 2.2 Parameter selection
2.3 Computational complexity
Each iteration requires $k+1$ evaluations of $F$, forms $\mathbf{B}^{\top}\mathbf{B}$ and $\mathbf{B}^{\top}\mathbf{r}$ in $O(mk^{2})$ time, solves a $k\times k$ system in $O(k^{3})$ time, and computes $\mathbf{P}\mathbf{w}$ in $O(nk)$ time. Storage for $\mathbf{P}$ and $\mathbf{B}$ costs $O(nk+mk)$ memory, while the Gram factorization adds $O(k^{2})$, yielding total per-iteration memory of $O(nk+mk+k^{2})$. The per-iteration computational cost is therefore $O((k+1)C_{F}+mk^{2}+k^{3}+nk)$ for residual cost $C_{F}$. When residual evaluations dominate, the linear algebra overhead remains negligible.
@@ 3 Numerical experiments
All experiments use the parameter choices described in section 2.2. Each solver receives 7,5007{,}500 residual evaluations and a maximum runtime of $30$ seconds. The metrics recorded include the best-so-far residual norm, the corresponding parameter error $\lVert\mathbf{x}-\mathbf{x}_{\star}\rVert_{2}$, and cumulative runtime. RSES sets $k$ from eq. 11 and maintains a fixed probe scale $\sigma$.
Results average ten independent trials per benchmark. The comparisons pit RSES against NEWUOA [powell2006newuoa], xNES [wierstra2014natural], Gauss-Newton (GN) with finite-difference Jacobians, ensemble Kalman inversion (EKI) [iglesias2013ensemble], and Adam [kingma2014adam] on benchmarks where gradients are available. Tables report mean residual, parameter error, and elapsed time at a shared stopping index: the first evaluation where the mean RSES distance falls within $1\%$ of its minimum. Every solver reports its values at this index, enabling fair comparison at a matched stopping point rather than at arbitrary terminal iterates. Convergence curves show mean best-so-far distance with root-mean-square bands on a shared evaluation axis.
The benchmarks escalate in difficulty, covering: a $2\times 2$ linear system, low-dimensional stochastic Brownian calibration, noisy multilayer-perceptron regression with up to 4,3534{,}353 parameters, and nonlinear deconvolution with $128$ parameters under both intact and perturbed residual weightings.
@@ 3.1 Linear algebraic system
The first test solves the $2\times 2$ linear system
@@ 3.2 Brownian drift and diffusion
and the goal is to tune the parameters
@@
$\hat{m}(\boldsymbol{\theta})=\tfrac{1}{N}\sum_{j=1}^{N}X_{1}^{(j)},\qquad\hat{v}(\boldsymbol{\theta})=\tfrac{1}{N}\sum_{j=1}^{N}\bigl(X_{1}^{(j)}-\hat{m}(\boldsymbol{\theta})\bigr)^{2}$ (16)
@@
Each trajectory evolves via Euler discretisation with $32$ steps of size $1/32$, averaged over 4,0964{,}096 paths.
RSES tracks the target despite Monte Carlo noise and achieves the smallest residual and parameter error. EKI follows closely in accuracy but runs slower. NEWUOA converges fastest yet stops far from the target. xNES reaches moderate accuracy with cost similar to RSES, while GN overshoots because noisy finite differences distort the Jacobian approximation (tables 2 and 3).
@@ 3.3 Multilayer perceptron regression
A one-dimensional regression task probes performance on a nonconvex landscape. The target signal follows
$y(x)=\sin(3x)+0.3x$ (18)
with additive Gaussian noise of standard deviation $0.05$. A two-hidden-layer multilayer perceptron with widths $8$, $16$, $32$, and $64$ fits these data using the smooth activation
$\phi(s)=\frac{s}{\sqrt{1+s^{2}}},$ (19)
which belongs to the softsign family and keeps derivatives bounded [glorot2010understanding], together with a Tikhonov penalty of $10^{-6}$.
With scalar input $x$, the network computes
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{1}\times 1}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{2}\times d_{1}}$, $\mathbf{W}_{3}\in\mathbb{R}^{1\times d_{2}}$, and $(d_{1},d_{2})$ takes one of the four width pairs. The Tikhonov residual
$F(\boldsymbol{\theta})=\begin{bmatrix}\hat{y}_{\boldsymbol{\theta}}(x)-y\\ \sqrt{\lambda/2}\,\boldsymbol{\theta}\end{bmatrix}$ (21)
stacks data misfit and $\ell_{2}$ regularisation, and the loss reads
@@
This surface contains broad plateaus and narrow valleys because the activation saturates and the regulariser couples all parameters, so gradient methods can stall even on this small dataset.
RSES uses $k=20$ (following $k=4+\lfloor 3\log m\rfloor$ with $m=256$ data points; the Tikhonov terms do not contribute to $m$ since the regularisation adds no problem complexity), $\sigma=0.01$, ridge scale $\beta=10^{-5}$, and approximately 20,00020{,}000 residual evaluations spread across $952$ iterations. Each iteration evaluates Gaussian probes of the Tikhonov residual without forming gradients. Adam [kingma2014adam] receives the same 20,00020{,}000 gradient evaluations. The comparison omits xNES, EKI, and GN on this task because their sampling and covariance costs scale poorly with these parameter counts. NEWUOA uses the same evaluation budget but slows sharply, exceeding the $30$-second time limit once the width reaches $16$.
Table 3 reports final losses $\mathcal{L}(\boldsymbol{\theta})$ and wall-clock time. RSES attains the lowest loss at every width while running faster than Adam and orders of magnitude faster than NEWUOA on wider networks. Adam remains competitive in loss but slower in time. NEWUOA stalls at substantially higher loss once the width reaches $16$.
Figure 4 shows residual histories across widths and the resulting best fits per algorithm. RSES descends steadily while Adam plateaus before recovering. The RSES fits track the noisy target closely without overfitting. These results indicate that RSES navigates the nonconvex MLP landscape reliably at this scale, though larger data regimes will require batching and may still favour tuned gradient-based training.